OpenStudy (anonymous):

If i =sqrt -1 and if x = 5 – 4i, then 1/x^2 -1=

OpenStudy (phi):

is this \[ \frac{1}{x^2 -1} \] ?

OpenStudy (anonymous):

We have to find \(x^2\) first..

OpenStudy (anonymous):

@rosebird don't keep yourself that busy. :P

OpenStudy (anonymous):

yeh its the same as u write @phi

OpenStudy (anonymous):

Yeah I am here only. :P

OpenStudy (anonymous):

what is the solution?

OpenStudy (anonymous):

\[x^2 -1 = (x-1)(x+1)\]

OpenStudy (anonymous):

Place value of \(x\) there in the brackets.

OpenStudy (anonymous):

solve it

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

1/41-40i

OpenStudy (anonymous):

Wait. :P

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Give me some time to solve it. :P

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

Check it once again: \[(4-4i)(6-4i) = ?\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

What is \((x-1)\) ??

OpenStudy (anonymous):



OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Getting?

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

So what it is: \[(4-4i)(6-4i)\]

OpenStudy (anonymous):

\(-40i\) is good, but check for \(41\).. :(

OpenStudy (anonymous):

answer?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

you worry much for answers, we are half way, or less than that yet. :P

OpenStudy (anonymous):

the options are 5-3i 4+i 1+5i/208 1-5i/208 @waterineyes

OpenStudy (anonymous):

You said : \[\frac{1}{41 - 40i}\] I said, \(-40i\) is correct, but check for \(41\) again, how you got it?

OpenStudy (anonymous):

You have two options: 1. Do whatever I am trying to say. Or 2. Follow the first one.. Which one you like? :P

OpenStudy (anonymous):

what are u saying

OpenStudy (anonymous):

We have not reached answer yet, we will be there but we have time now to reach there? Why you worry for answers?

OpenStudy (anonymous):

then post ur answer and solution here i'll be back

OpenStudy (anonymous):

Simply multiply it correctly: \[(4-4i)(6-4i)\]

OpenStudy (anonymous):

Sorry I can't do that. :)

OpenStudy (anonymous):

after multiplying its 1/40-52i

OpenStudy (anonymous):

Haven't you study our Code of Conduct yet? If you have not then read this, click here: \(\href{http:///openstudy.com/code-of-conduct}{\tt{Code \ Of \ Conduct.}}\)

OpenStudy (anonymous):

@sirajawahid

OpenStudy (anonymous):

Please do the calculations carefully.

OpenStudy (anonymous):

you are multiplying it wrongly. :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Probably it is happening because you want to go with speed of light. :P

OpenStudy (anonymous):

Calm down, do the calculations carefully, you have got all the time to learn. :)

OpenStudy (anonymous):

@waterineyes, explain it to me please :)

OpenStudy (anonymous):

From where? Till where you got it?